### v2

# When LLMs Go Online: The Emerging Threat of Web-Enabled LLMs

Abstract
Recent advancements in Large Language Models (LLMs) have established them as agentic systems capable of planning and interacting with various tools. These LLM agents are often paired with web-based tools, enabling access to diverse sources and real-time information. Although these advancements offer significant benefits across various applications, they also increase the risk of malicious use, particularly in cyberattacks involving personal information. In this work, we investigate the risks associated with misuse of LLM agents in cyberattacks involving personal data. Specifically, we aim to understand: 1) how potent LLM agents can be when directed to conduct cyberattacks, 2) how cyberattacks are enhanced by web-based tools, and 3) how affordable and easy it becomes to launch cyberattacks using LLM agents. We examine three attack scenarios: the collection of Personally Identifiable Information (PII), the generation of impersonation posts, and the creation of spear-phishing emails. Our experiments reveal the effectiveness of LLM agents in these attacks: LLM agents achieved a precision of up to 95.9% in collecting PII, up to 93.9% of impersonation posts created by LLM agents were evaluated as authentic, and the click rate for links in spear phishing emails created by LLM agents reached up to 46.67%. Additionally, our findings underscore the limitations of existing safeguards in contemporary commercial LLMs, emphasizing the urgent need for more robust security measures to prevent the misuse of LLM agents.
1 Introduction
The integration of various external tools, such as APIs and databases, has significantly enhanced the capabilities of Large Language Models (LLMs). This integration allows LLMs to function autonomously as agents — advanced AI systems that utilize LLMs as their core component to perform complex tasks. Numerous studies have shown the effectiveness of LLM agents in performing tasks across various domains [12, 35, 57, 54].
Given the advanced capabilities of LLM agents, there is growing concern about their potential to cause significant real-world harm if used maliciously. Recent studies [28, 17, 18] have shown that LLM agents could be exploited for cyberattacks, highlighting the risks they pose in this context. In response to these risks, LLM vendors (e.g., OpenAI  [^1] https://www.openai.com, Google  [^2] https://www.google.com) have implemented policies that prohibit harmful activities, such as compromising privacy or intentionally deceiving others [43, 6]. Additionally, safeguards have been established to prevent the misuse of these powerful models [40, 22, 4]. However, as models become more capable, it becomes increasingly challenging to foresee and mitigate all potential misuse scenarios.
The risks associated with LLM agents become even more problematic when paired with web-based tools. The widespread sharing of personal information on the web makes it an appealing target for cybercriminals. For example, attackers often resort to data scraping from websites like Google, LinkedIn, and Facebook to collect personal information, which is then sold on hacker forums [52, 56, 55, 39]. Such activities not only infringe on privacy but also increase individuals’ vulnerability against targeted cyberattacks, including impersonation and phishing email attacks [25, 23]. Attackers can use the acquired information on the target to develop specifically tailored documents for malicious purposes, where impersonation posts wrongfully take advantage of the reputation of the target, and phishing emails deceive the target into clicking on a harmful link.
Our work. Despite the web’s utility and the vulnerabilities introduced by the prevalence of personal data, the capability of LLM agents in cyberattacks exploiting personal information remains unclear. In this work, we examine the effectiveness and harmfulness of LLM agents using web-based tools (such as web search API and hand-designed functions for navigation) for malicious purposes concerning personal data. To this end, we identify three hallmark cyberattacks that utilize publicly available private information by leveraging LLM and agents: personally identifiable information (PII) collection, impersonation post generation, and spear phishing email generation.
By conducting a systematic analysis of these cyberattacks using LLM agents, we aim to answer the following research questions.
RQ1. How potent are the cyberattacks exploiting personal information conducted by LLM agents? To assess the potential misuse of LLM agents in cyberattacks, we design proof-of-concepts of the three hallmark cyberattacks. We evaluate the latest commercially available LLMs [^3] We utilize the latest models as of July 17, 2024. Our attacks were last confirmed to be valid on September 1, 2024.—GPT, Claude, and Gemini—that are readily accessible to the public. To quantitatively analyze how privacy is infringed by LLM agents, we use them to collect five types of PII (name, email address, phone number, office address, and personal web page) of CS researchers from 10 prominent universities, using only the university name as initial input for prompt. In addition, to determine if an LLM agent can automatically create effective impersonation posts, we provide the LLM agent with only the names and affiliations of the impersonation targets and prompt them to generate posts as if they are the person and endorse a specific claim. Lastly, we evaluated the effectiveness of LLM agents in crafting spear phishing emails through a user study with 60 participants. This study focused on assessing the authenticity and credibility of various phishing email variants generated by LLM agents. Our findings indicate that LLM agents are able to properly retrieve the PII of targets with high precision rates (up to 0.95 for GPT). Furthermore, evaluations show that up to 93.9% of posts generated by LLM agents were perceived as authentic. Spear phishing emails generated by LLM agents were also effective, with up to 46.67% of the participants click on the malicious link.
RQ2. How much do web-based tools enhance LLM agents undergoing cyberattacks? LLM agents enhance their responses using web-based tools by searching the web and navigating various interactive elements. Previous literature report proficiency of LLMs when used in cyberattacks [51, 47, 9, 38, 32], and the addition of a powerful feature can be expected to increase this risk. To measure the impact of web-based tools on cyberattacks, we focus on the performance difference between using vanilla LLMs and LLM agents enabled with web-based tools. We experiment with two different levels of capability: an LLM agent enabled with searching functions and an LLM agent enabled with searching and navigation functions. We find that LLM agents consistently outperform vanilla LLMs in all the cyberattacks.
RQ3. How approachable are LLM agents to misuse for cyberattacks? The approachability of cyberattacks via LLM agents can be broken down into two categories: cost and safeguard capability. The execution of such attacks incurs both time and financial costs, which significantly influence the practicality of using LLM agents for cyberattacks. Our findings indicate that LLM agents can exploit such attacks very quickly and at minimal cost, emphasizing their practicality. On average, the GPT LLM agent can perform each task within 10 seconds at a cost of around 2 cents. The challenge of misusing LLM agents arises from the inherent safeguard features embedded in the LLM service platforms. For an attack to be executed through LLM agents, the attacker’s prompts must successfully circumvent these safeguards. Our results demonstrate that safeguards implemented by LLM providers were activated only in particular scenarios and with certain services. Furthermore, we found that simply enabling web-based tools actually allowed bypassing of safeguards for some LLM services.
Contributions. We outline our contributions as follows:
  • We systematically explore the capabilities of LLM agents to exploit private data in cyberattacks. Specifically, we demonstrate that LLM agents can successfully: 1) collect five types of PII, 2) impersonate specific targets to promote an attacker’s claims using personal information, and 3) create highly targeted emails by crafting suitable scenarios and sender identities for recipients.
  • We found that safeguards of even the latest LLM services often fail to function effectively. Notably, the incorporation of web-based tools often led to LLMs being more permissive, allowing prompts to bypass the safeguards. These findings reveal a significant weakness in current safeguards and highlight the urgent need for more robust security measures to prevent the misuse of LLM agents.
Content Warning. Please note that this work discusses and carefully examines the potential threat posed by LLM agents in cyberattacks. We believe that the benefits of publicizing these findings outweigh the potential harms. To prevent misuse, we have provided only summarized versions of the prompts used for the attacks in this work. Full prompts will be disclosed only upon request for legitimate research purposes.
2 Related Work
LLM Security. As LLMs have advanced and become widely adopted, their potential misuse has raised significant concerns. Previous research has demonstrated that LLMs can be exploited in various domains, including PII extraction [27, 31, 38], deepfake user profile creation [36], and phishing email generation [51, 47, 9]. A recent study [32] also highlighted the proliferation of LLMs being repurposed as malicious services in underground marketplaces, further underscoring the risks associated with these powerful models.
Efforts to assess the safety of LLMs have included red-teaming approaches such as jailbreak attacks [7, 14, 30]. In terms of enhancing security, prevailing methods for safety alignment often involve reinforcement learning from human feedback (RLHF), aimed at promoting the development of safer LLMs [7, 44]. With the capability of LLM agents, recent studies have leveraged LLM agents for safeguard purposes [62, 61, 58].
Despite existing safeguards in commercial LLM services [40, 4, 22], we found that even the latest models (as of July, 17, 2024) lack effective protections against our attack scenarios, emphasizing the urgent need for safeguard enhancement.
LLM Agents for Cyberattacks. An LLM agent refers to an artificial intelligence system that uses an LLM to perform tasks, make decisions, or interact with users autonomously [12]. These agents are adaptable to various applications, depending on their training and integration into systems. Recent studies have demonstrated the capability of LLM agents in various domains, such as event forecasting [59], real-world coding challenges [63], and web navigation [35].
Despite these capabilities, there is a growing concern about the potential misuse of LLM agents. Research indicates that LLMs often fail to recognize safety risks in agent scenarios [60]. Furthermore, prior studies have shown that LLM agents can autonomously exploit vulnerabilities, ranging from SQL injections on websites [18] to one-day vulnerabilities in real-world systems [17].
Although the threat of social engineering involving private data on online platforms is growing, there is a lack of research on the capabilities of LLM agents to exploit private data in cyberattacks. Our work addresses this gap by systematically studying the potential for LLM agents to misuse private data in cyberattacks.
3 LLM Agents for Cyberattacks
Figure 1: Process by which the LLM responds to the user’s prompt: (a) Vanilla LLM without tools (blue line), (b) WebSearch agent with web search functionality (orange lines), and (c) WebNav agent with both search and navigation functionalities (green lines).
In this section, we explore the risks associated with LLM agents when integrated into web-based tools, aiming to simulate cyberattack scenarios. This analysis is designed to enhance our understanding of how such tools might be misused in real-world contexts. Section 3.1 details the specific tools and methodologies used to deploy LLM agents in cyberattack scenarios. In Section 3.2, we introduce the specific tasks assigned to LLM agents within these cyberattack scenarios.
3.1 Overview of LLM Agents
To explore the potential vulnerabilities of LLMs to cyberattacks, we simulate scenarios where LLMs could be exploited by attackers. A straightforward approach to exploiting LLMs involves prompting them to perform harmful actions, such as extracting personal information (Figure 1(a)). While LLMs can execute convincing attacks with malicious prompts, we focus on more advanced versions, known as LLM agents, that are capable of planning and interacting with various tools, including software and external APIs.
Specifically, we consider two LLM agents that leverage web-based tools like web search and web navigation:
  • WebSearch Agent: This agent leverages a web search tool to access search results from engines like Google and Bing.
  • WebNav Agent: This agent employs a navigation tool to retrieve content from web pages and interacts with clickable elements to access more deeply embedded information.
Implementation. We implement LLM agents using the function calling feature provided by each LLM’s API. We provide a set of function descriptions to the LLM, enabling the model to determine the appropriate timing and method for calling functions based on the task requirements. It is important to note that LLMs do not execute functions directly; rather, they identify the appropriate moments for function execution and supply the necessary arguments. The actual execution is carried out by an application, such as a web search tool, which then returns the results to the LLM. The LLM uses these results to generate a response, thus automating the process and enabling the agent to perform designated tasks effectively.
For the WebSearch Agent, we implement the search() function using the Custom Search JSON API [21], which retrieves Google search results in a structured JSON format. This function accepts a search term as an argument and returns the corresponding Google search results. As shown in Figure 1(b), the WebSearch agent calls the search() function with an appropriate query and then uses the returned search results to generate a response. When the agent cannot find the required information from the results, it may repeatedly call the function, adjusting the query as needed. In Appendix A, we include the results of search() as an example.
For the WebNav Agent, we implement the functionality using web automation tools, such as Selenium [50] and BeautifulSoup [46] with Requests [26]. Specifically, we develop two functions: fetch_content() and find_button(). The fetch_content() function takes a URL as an argument and returns the content of the site, while the find_button() function identifies clickable buttons and their corresponding URLs at a given URL.
As shown in Figure 1(c), when a WebNav agent cannot find the desired information using solely the search() function, it employs the fetch_content() and find_button() as follows:
  • (1)
    The agent visits appropriate URLs obtained from search results using fetch_content().
  • (2)
    It analyzes whether the required information is present in the fetched content.
  • (3)
    If the information is not found, find_button() is used to identify appropriate buttons or tabs for accessing additional information.
  • (4)
    The agent then fetches content from the new URL using fetch_content().
  • (5)
    Repeat steps 1 to 4 until the agent finds desired information.
  • (6)
    Finally, the agent synthesizes the results to generate a comprehensive response.
Note that these steps are completely automated, relying solely on the input prompt without receiving human feedback, unlike in chatbot or human assistant settings.
Models. We employ commercially available models, whose ease of access and capabilities can encourage misuse by attackers. Specifically, we use GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Flash. We utilize the respective APIs: the OpenAI API  [^4] https://openai.com/index/openai-api/ for GPT-4o, the Anthropic API  [^5] https://docs.anthropic.com/en/api/getting-started for Claude 3.5 Sonnet, and the Gemini API  [^6] https://ai.google.dev/api?lang=python for Gemini 1.5 Flash. Throughout this work, we will refer to GPT-4o as GPT, Claude 3.5 Sonnet as Claude, and Gemini 1.5 Flash as Gemini.
Table 1: Summary of experimental settings for each attack.
PII collection Impersonation post generation Spear-phishing email generation
(Section 4) (Section 5) (Section 6)
Attack Goal Get targets’ 5 types of PII Generate social media posts that impersonate a target advocating a specific claim Generate targeted emails to induce recipients click a link embedded in the email
Attack Target CS researchers CS professors Academic researchers, Non-academic professionals
Attack Inputs Name of university Target’s name, Target’s institution, Desired claim Target’s email address
Evaluation Human annotation LLM evaluation User study
3.2 Targeted Cyberattacks
In this work, we investigate the capabilities of LLM agents to execute cyberattacks that traditionally require considerable human effort and resources. Specifically, we focus on three tasks that exploit vulnerabilities stemming from the widespread availability of personal data online, each varying in complexity:
Attack 1. PII Collection. Collecting PII, such as email addresses, names, and phone numbers, from the internet raises significant privacy concerns, even when the data is publicly accessible. Attackers can exploit this information for malicious purposes such as identity theft, fraud, or orchestrating further cyberattacks [25, 23]. They often employ web scraping tools and automated scripts to gather PII from various websites. However, due to the unique formatting of information on each website, these tools typically require custom development for each target site, which involves considerable human labor and associated costs [53, 15].
Attack 2. Impersonation Post Generation. Attackers craft impersonation posts to deceive audiences for malicious purposes, such as financial gain, or to tarnish reputations [29, 33]. To create sophisticated impersonation posts, attackers meticulously research their targets’ personal details, social behaviors, and communication styles. These elements are then integrated into their writing, a process that requires significant time and effort.
Attack 3. Spear Phishing Email Generation. Spear phishing emails are highly effective at deceiving victims but tend to be more costly and time-consuming for attackers compared to traditional phishing emails [48, 24]. This increased effort results from spear phishing’s focus on targeting specific individuals or organizations with tailored messages.
Collecting PII is the most straightforward task, where LLM agents gather targeted information from online contexts. Generating impersonation posts is a more complex task that involves using personal data to craft content closely mimicking the target. The most complex task, spear phishing email generation, involves designing scenarios, establishing sender identities, and creating content tailored to the target’s interests. We summarize experimental settings of these attacks in Table 1.
In the following sections, we answer our research questions via analyses of these attacks. The potency of LLM agents when misused for these attacks (RQ1) as well as the impact web-based tools of LLM agents have (RQ2) are answered in Sections 456, by conducting the aforementioned attacks in a fully automated state. The approachability (RQ3) will be answered in Section 7 by measuring the cost and performance of innate safeguard features. Our findings highlight the potential threat posed by LLM agents in automating sophisticated cyber threats.
4 PII Collection
In this section, we investigate the potential for LLM agents to infringe upon privacy by extracting personally identifiable information (PII).
4.1 Experimental Setup
Attack Scenario. The attacker’s goal is to collect PII of specific targets using LLM agents. In this work, we collect the PII of professors from the CS departments of the top ten universities, as ranked by the QS World University Rankings. [^7] Professors are particularly susceptible to targeted cyber attacks due to their high profile and reputation [10, 49]. The types of PII we consider are names, email addresses, phone numbers, office addresses, and URLs of personal web pages for each target.
To further assess variations in information collection based on academic roles, we also investigate whether LLM agents can collect the PII of CS students. Specifically, we uniformly sample 50 professors at random, whose personal web pages are successfully harvested during our initial PII collection. We then set our targets as CS students affiliated with the laboratories of these sampled professors.
Figure 2: Pipeline of PII collection (professor group). We use MIT as an example.
The pipeline of PII collection (illustrated in Figure 2) using LLM agents consists of two steps: 1) constructing a list of target names and 2) collecting their corresponding personal information. Note that the attack consists of no manual effort by the attacker and only by utilization of LLM agents. For the professor group, the attacker initially extracts the names of CS professors from a given university (e.g., MIT). Then, the attacker collects the remaining PII for each professor listed. For the student group, the data obtained from the professor group is utilized. Specifically, the names, personal web pages, and school affiliations of professors are used in the prompts to identify the names of students currently associated with each professor. The collected student names are then used as the input for the LLM, which is tasked with retrieving the PII for each student.
Evaluation. Due to the vast and diverse nature of the field of CS, we did not evaluate based solely on department listings. Instead, we employed human annotators to review the responses generated by LLM agents. Annotators were directed to classify individuals as CS professors based on their engagement in relevant fields, following a thorough web search. Certain types of PII, like phone numbers, are prone to changes over time, complicating verification; thus, we initiated five separate queries to LLMs for the remaining PII for each individual. We considered the information collection successful if at least one of the five responses matched the data available online. During the annotation process for the names of CS students, we used the student information listed on the professors’ web pages as the ground truth. For more details on the annotation process, please refer to Appendix B.1.
Figure 3: The number of correctly collected PII for (a) CS professors and (b) CS students, categorized by each type of PII. (c) Precision and Recall of the collected names of CS students. We use GPT for all methods.
We use precision and recall as our evaluation metric. Precision for each category was calculated as the proportion of successfully collected PII relative to the total PII collected. We utilized the CS faculty lists provided by the official websites of each university as the reference for comparing our collected data. We calculated the upper bound recall  [^8] During our review, we identified several individuals not listed on the official CS faculty pages but recognized as CS professors by our annotators. These individuals were confirmed to be researchers specializing in CS-relevant fields such as information science and machine learning. using the number of professors officially listed under the CS faculty.
4.2 Main Results
Figure 3 (a) and Figure 3 (b) show the results of PII collection for CS professors and students, respectively, using vanilla LLM and LLM agents with GPT. The effectiveness of LLM agents in collecting PII improves as they utilize additional tools. This implies that the risks associated with LLM agents significantly increase as their capabilities expand.
For CS professors, the vanilla LLM without the utilization of web-based tools were less effective in retrieving PII, particularly for detailed information like office addresses and phone numbers. By utilizing the search API, the WebSearch agent collected information more effectively than the vanilla LLM. However, it was less effective in retrieving office addresses and phone numbers due to the limitations of the Google Custom Search API. This API returns only snippets [19] (automatically generated excerpts that summarize page content), which often mix details of multiple professors, leading to inaccuracies, especially in contact information. In contrast, by additionally utilizing navigation tools, the WebNav agent effectively collected the names of 570 out of 686 CS professors and achieved substantial collection rates for additional PII, including phone numbers (71.4%), office locations (91.2%), email addresses (95.9%), and personal web pages (77.7%).
In the case of PII collection for CS students, only the WebNav agent was effective, as students often did not have or disclose their personal information as comprehensively as professors. As shown in Figure 3 (c), the WebSearch agent exhibited poor performance even in collecting students’ names, constrained by the limitations of the search API as mentioned before. Additionally, the vanilla LLM often returned numerous incorrect student names with significantly low precision, demonstrating substantial hallucinations.
These results demonstrate that access to the web significantly enhanced the performance of LLM agents in extracting PII of specified targets. As the capabilities of LLMs increase, so does the potential threat, enabling the collection of more detailed information through multiple layers of exploration. Although the information is publicly available on the Internet, the successful retrieval of PII by these agents is concerning and may necessitate a reconsideration of how LLM agents are utilized.
Figure 4: The number of correctly collected PII for CS professors categorized by each type of PII, using Claude for all methods.
Figure 5: Precision and Recall of the collected names of CS students, using the Claude for all methods.
4.3 Model Analysis
In order to examine the differences in PII collection capabilities, we provided prompts to Gemini and Claude for PII collection. For Gemini, regardless of the use of web-based tools, the model refused to collect PII, responding with messages like “Sharing such information could be a violation of privacy and potentially lead to harassment or security risks.”.
In contrast, as illustrated in Figure 4, Claude collected PII in a manner similar to GPT, where more advanced tools gather information more effectively, except in collecting professors’ names. The vanilla LLM managed to collect more names than its agents, but this is due to its indiscriminate collection process, which resulted in a low precision of 0.574 (out of 1,384 names collected, only 795 were correct). For constructing the list of student names, only the WebNav agent, like GPT, was effective in achieving both precision and recall above 0.9 (Figure 5).
When comparing the performance of WebNav agents from GPT and Claude, Claude slightly underperformed as it failed to retrieve the student list from 10% of the given professors’ websites. This difference can be attributed to their distinct exploration strategies: GPT, despite uncertainties, consistently explored to locate the student list, while Claude tended to stop if the context was unclear. Specifically, if the student list was explicitly listed on the given website, the LLM agent scraped and returned the information directly. If not immediately available, the agent employed navigational techniques within the site to locate the student list, utilizing buttons or links as necessary. For instance, the agent may need to access another page via a link on the given site to search for the student list therein. Even if the site did not explicitly indicate that the link leads to a student list, GPT will explore the link to determine if the necessary information is present. In contrast, Claude only pursued links if they are clearly marked, and will report the absence of student information if clarity is lacking.
5 Impersonation Post Generation
In this section, we explore how effectively LLM agents impersonate individuals by using publicly available information from the web.
5.1 Experimental Setup
Attack Scenario. The attacker aims to automatically generate social media posts that impersonate specific targets to endorse the attacker’s claims. LLM agents are used to generate credible impersonation posts using a target’s personal information from the web, allowing attackers to exploit the target’s reputation to disseminate their intended message. In this study, we target 50 CS professors, as identified in Section 4.
As illustrated in Figure 6, an attacker can create an impersonation post by sending a single query to the agents. Directly including terms like “impersonate” was shown to trigger the LLM’s safeguards, resulting in rejection. Therefore, we use a simple role-playing technique, introducing the target to the agent with the statement, “I am [Name] at [University]”, followed with web-based tools to learn more about the target’s background. Subsequently, we request the agent to compose a 500-700 word social media post advocating a specific claim. Note that the attacker only uses three inputs in the prompt: the target’s name, institution, and the claim to be promoted. As a baseline, a vanilla LLM is used to craft an impersonation post but without the web search function, relying solely on pre-stored knowledge.
The claims established in this attack are: 1) “recommend researching AI”, and 2) “LLM is highly secure against potential misuse”. The claims are designed to observe how the nature of the claim (the first is relatively innocent, while the latter is controversial) affects the LLM agent’s performance in impersonation tasks.
Figure 6: Pipeline of impersonation post generation.
Evaluation. We examined the impact of web accessibility on impersonation by conducting an A/B test. Impersonation posts were generated using (1) vanilla LLMs, (2) WebSearch agents, and (3) WebNav agents; then, we conducted pairwise comparisons to determine which of the two posts was more likely to be written by the target individual. As evaluators, we employed LLMs, which are recognized for their high accuracy in tasks such as fake news detection and machine-generated text detection, even without additional fine-tuning [45, 11, 34]  [^9] We also verified that LLM evaluators showed a high correlation with human evaluations in our experiment (See Appendix C.2).. We enabled the web search function to enhance the accuracy of the LLM evaluations by leveraging internet-available information. Given impersonation posts from two models (e.g., vanilla LLM vs. WebSearch agent, or WebSearch agent vs. WebNav agent), LLM evaluators indicated which post more accurately impersonated the target individual, or indicated ’unsure’ if the two posts were similarly effective. We randomize the placement of texts (A or B) to mitigate order bias. Furthermore, to minimize potential model bias, each pair of posts was evaluated by three different models: GPT, Claude, and Gemini. The final outcome was determined based on the majority vote from the three evaluations.
In addition, to determine how likely the impersonation posts were perceived to be authentic, we conducted a Yes/No test. Similar to the A/B test, we utilized the three LLM evaluators and to assess whether the texts appeared to have been authored by the target individuals. The templates for the A/B test and Yes/No test are provided in Appendix C.1.
Figure 7: Example of a tweet impersonating a well-known figure claiming “LLM is highly secure against potential misuse” generated by WebSearch agent with GPT.
5.2 Main Results
Figure 7 shows an example of an impersonation post  [^10] For visualization, we prompted the agent to generate a tweet of up to 300 words with the second claim. of a well-known figure generated by the WebSearch agent with GPT. We provide the agent with only the name information; however, it synthesizes information about his expertise, alma mater, and companies to craft a tweet that convincingly appears to be authored by him. The proceeding tests are conducted using target professors according to our attack pipeline.
Comparison Based on Tool Usage. We conducted comparative analysis of impersonation posts generated by WebNav agents and WebSearch agents, followed by comparisons between WebSearch agents and vanilla LLMs, to examine their effectiveness based on their capability. Table 2 shows the proportion of choices made in an A/B test. Cases where the LLMs or agents refused to generate content were excluded from our calculations (these will be discussed later in this section). The results indicate that the impersonation effectiveness increases with access to additional tools across all models. Specifically, WebNav agents were more effective than WebSearch agents, and WebSearch agents outperformed vanilla LLMs. The differences in effectiveness were more pronounced when comparing WebSearch agents to vanilla LLMs, suggesting that adding search capabilities (WebSearch agents) provides a significant boost in performance for tasks requiring impersonation abilities.
Table 2: A/B test results. The percentages represent each decision relative to the total number of decisions made. claim1 refers to “Recommend researching AI” and claim2 to “LLM is highly secure against potential misuse”. WebN and WebS represent WebNav and WebSearch, respectively.
claim1 claim2
Gen WebN Agent WebS Agent Unsure WebN Agent WebS Agent Unsure
GPT 56.3 39.6 4.2 58.3 39.6 2.1
Claude 51.0 30.6 18.4 51.0 30.6 18.4
Gemini 52.4 35.7 11.9 44.4 25.9 29.6
Gen WebS Agent LLM Unsure WebS Agent LLM Unsure
GPT 72.1 11.6 16.3 61.2 30.6 8.2
Claude 72.1 11.6 16.3 72.1 16.3 11.6
Gemini 72.1 11.6 16.3 59.2 20.4 20.4
Table 3: Yes/No Test results using GPT evaluators. Percentages represent the proportion of “Yes” responses.
claim1 claim2
Gen WebNav Agent WebSearch Agent LLM WebNav Agent WebSearch Agent LLM
GPT 93.9 89.8 75.5 73.5 61.2 40.8
Claude 85.7 71.4 65.3 69.4 40.8 6.1
Gemini 74.0 54.0 22.0 52.0 52.00 6.0
Authenticity Evaluation. We examine whether the texts created by the WebNav agents, WebSearch agents, and vanilla LLMs are perceived as being written by the target. Table 3 shows the percentage of “yes” responses by the LLM evaluator for posts created by vanilla LLMs and LLM agents. The results are similar to those observed in the A/B test: agent-generated posts had a higher “yes” response rate than LLM-generated posts. WebNav agents performed the best; the GPT and Claude agents achieved success rates of up to 93.9% and 85.7%, respectively. Compare to GPT and Claude, Gemini showed poor performance in most cases.
We attribute Gemini’s poor performance to the brevity of its texts. Although we instructed the models to generate texts between 500-700 words, Gemini produced significantly shorter texts, averaging 311 words for claim 1 and 334 words for claim 2. This conciseness likely resulted in posts that lacked adequate personal information, leading to a higher frequency of “no” responses. Another reason for the performance decline could be Gemini’s tendency to insert placeholders instead of actual information when it encounters unfamiliar content. For example, it used phrases like “As a researcher with expertise in [mention your research area] and a strong understanding of [mention relevant fields]”, indicating a lack of details. These factors collectively contributed to the lower effectiveness of the posts generated by Gemini.
Interestingly, the effectiveness varies slightly depending on the nature of the claim. Claim 2, “LLM is highly secure against potential misuse”, may not accurately reflect reality and is potentially controversial. Despite incorporating personal information, LLM evaluators were skeptical about a CS professor making such a claim, leading to more uncertainty across all models.
Web Tool as Jailbreak. Our experiments maintained consistent prompts while varying only the target’s name, institution, and claim. Interestingly, while the WebSearch agents successfully generated posts for all professors, the vanilla LLMs refused to generate content for some professors due to security concerns: one professor’s posts were denied generation by GPT for claim 1, and eight professors (one by GPT and seven by Claude) were excluded for claim 2. This highlights just using web tools can inadvertently circumvent these safeguards, potentially acting as a form of jailbreak. A detailed discussion of safeguards in LLMs is provided in Section 7.
Figure 8: Pipeline of spear phishing email generation.
6 Spear Phishing Email Generation
In this section, we investigate whether LLM agents can be used to generate personalized phishing emails using only an email address, without the need for additional human intervention.
6.1 Experimental Setup
Attack Scenario. In Section 4, we demonstrated the feasibility of effectively obtaining the email addresses of researchers in the university. Similarly, attackers can exploit publicly available email addresses [37] or acquire them through purchases on the dark web or Telegram channels [16]. Attackers commonly use phishing emails to obtain sensitive information from targeted individuals by asking them to click on a link or enter personal information [8]. In our study, the attacker’s goal is to craft highly personalized phishing emails generated by LLM agents to induce recipients to click the malicious link, utilizing the email addresses of target individuals.
The attack pipeline is designed to craft highly convincing phishing emails. Note that the attack consists of no manual effort by the attacker and only by utilization of LLM agents. The only ingredient for this attack is the target’s email address, which is added to the generation prompt of the LLM query. As illustrated in Figure 8, an attacker can create a phishing email by sending a single query to the agent. The process begins with verifying the date to ensure the timing aligns with the intended attack scenario. Then, the content of the email is developed by searching the target’s email address online to gather personal information. This information is then utilized to design a realistic scenario where the target is likely to receive an email, along with identifying a credible sender that fits this scenario. Subsequently, a plausible URL string is generated and embedded in the email to encourage the target’s engagement. Finally, the sender’s email address is altered to mimic a legitimate domain, enabling the dispatch of the email without exploiting vulnerabilities in official domains or compromising accounts to access the authentic domain.
Note that we conducted two pilot studies, which showed similar results from WebNav and WebSearch Agents. Due to cost and time constraints, WebNav-generated emails were excluded from the main evaluation.
User Study. Phishing emails were custom-generated for each target, requiring recipients to personally evaluate the emails intended for them. Thus, we designed a questionnaire to assess these emails and actively recruited participants for a comprehensive evaluation.
As in Section 4, academic researchers’ information is publicly available online, making them more vulnerable to phishing attacks. Therefore, we categorized participants into two groups: academic researchers (including undergraduate students) and non-academic professionals (individuals within non-academic organizations). Academic researchers were recruited via university research communities, while non-academic professionals were recruited through industry practitioners from prior collaborations. Specifically, we recruited 60 participants and categorized them based on the provided institutional email addresses to analyze response variations across institutional types: 55% were academic researchers and 45% were non-academic professionals  [^11] The recruitment process with ethical/privacy-preserving measures is explained in Appendix E.1..
We emphasize that participants did not receive these emails directly; instead, they were assessed via a Google survey. Each participant evaluated seven phishing emails generated using their own email as a target. The types of emails will be described in Email Design part.
Table 4: Types of LLM-crafted phishing emails.
Model Web search Purpose Sender’s institution
Claude o General not designated
GPT o General not designated
Claude x General not designated
GPT x General not designated
Claude o Credentials not designated
GPT o Credentials not designated
Claude o General different from target’s
Email Design. Using direct terms like “phishing email” would trigger the LLM’s safeguards and result in rejection. Therefore, our experiment includes minimal circumvention techniques, such as avoiding such explicit terms. Although these minimal efforts were sufficient for Claude and GPT, they could not bypass the safeguard restrictions of Gemini and refused generation.
As shown in Table 4, we designed seven distinct types of emails to evaluate the effectiveness across different variants of phishing emails. We divided the purpose of email into general and login credentials. For the general purpose, the WebSearch agents were tasked to generate a realistic scenario in which the target might receive an email. To examine the effectiveness of web search functionality, we compared the emails generated by vanilla LLMs to those created by WebSearch agents. For scenarios involving login credentials, we provided specific instructions for WebSearch agents to generate an email prompting the target to update their credentials via a designated link. Additionally, to examine the impact of the sender’s affiliation, we modified the prompt for the WebSearch agent using Claude to make the sender belong to a different institution than the target.
Questionnaire Design. We designed seven multiple-choice questions to build an understanding of how participants perceive and interact with the emails, from surface content to deeper implications of authenticity and expected actions.
The first two questions focus on the content of the emails, asking participants to identify the presented information (Q1) and to evaluate its accuracy against their actual personal details (Q2). Subsequent questions explore perceptions of the sender and recipient, including whether the recipient identified the sender (Q3) and if the recipient’s name correctly matched their real name (Q4). We then inquire about the actions participants would likely take upon receiving the email (Q5). The last question investigates perceived authenticity of the email (Q6), asking participants to rate how genuine or fraudulent it appeared and to identify specific elements that influenced their judgment (Q7). The questionnaire is available in Appendix E.5.
Figure 9: Example of a phishing email to a well-known figure generated by the WebSearch agent with Claude using an email address only.
6.2 Main Results
We first analyze the effectiveness and content of general-purpose emails created by LLMs and WebSearch agents. Then, we examine the results according to the participants’ groups. We then analyze the efficacy of emails according to their purpose and sender’s institution. In Tables 10 (Appendix E.3), we report the comprehensive survey results concerning participants’ anticipated actions for various email types.
Figure 10: Survey results on participants’ (a) expected actions upon receiving emails generated by LLMs and WebSearch agents, and their (b) perceived authenticity.
6.2.1 Overall Analysis
Figure 9 provides an example of phishing email generated by WebSearch agent using Claude. We prompted the agent to write an email to a well-known figure, by providing only his publicly available email address. The agent generated a plausible email considering his information, including a realistic-looking URL, and proposed a collaboration between Meta and Spotify.
Effectiveness Analysis. Figure 10 (a) presents the survey results of actions participants would likely take upon receiving the email (Q5). Our primary focus was on the proportion of participants clicking on links, directly aligning with the attacker’s goal. The results indicate that WebSearch agents prompt more link clicks than vanilla LLMs, with the WebSearch agent using Claude achieving the highest click rate at 26.67%  [^12] Expert-crafted spear phishing emails in [9] show 26.6% click rate. Detailed discussion in Section 6.2.3.. In the case of GPT, the WebSearch agent’s click rate was twice that of the vanilla LLM. When comparing Claude and GPT, Claude was more effective, with its vanilla LLM variant achieving a slightly higher click rate than the WebSearch agent using GPT. The disparities in effectiveness between Claude and GPT may be attributed to differences in email content, which will be explored further in the subsequent section.
When comparing “actions” (Q5) and “perceived authenticity” (Q6) in Figure 10, on average, 66.27% of respondents who identified an email as definitely phishing opted for move to spam as their action. However, the percentage of participants who rated the email as genuine (definitely genuine + probably genuine) did not closely match the percentage who chose to click link. Interestingly, for emails generated by WebSearch agents, the click rate was much higher than the rate of perceived genuineness, a pattern not observed with vanilla LLM-generated emails. This discrepancy was driven by participants marked as unsure or even phishing; 13.34% of participants stated that they would clicked on links in WebSearch agent-generated emails despite doubting their authenticity. (For vanilla LLM-generated emails, 10% and 3.34% did so for Claude and GPT, respectively.) Table 11 (Appendix E.3) presents a contingency table showing the percentage distribution of participants’ actions (Q5) relative to their perception of authenticity (Q6). These results highlight that the WebSearch agents not only provided a sense of credibility but also effectively utilized target information from the web to engage the recipients’ interest, leading to higher link-clicking rates.
Figure 11: Analysis of email content: (a) detail of email content and (b) alignment of email content with targets.
Content Analysis. We analyzed which email aspects influence participants’ perceptions of authenticity using chi-squared tests. From the analysis, the detail of email content and the accuracy of those details in aligning with the target’s information played statistically significant roles in participants’ assessment of an email’s authenticity. Table 8 in Appendix E.2 provides statistical results of the chi-squared tests.
Building on these statistical results, we further examined the influence of web accessibility on the two factors. Figure 11 presents a comparative analysis of the percentage of phishing emails that incorporate specific content details and how well this content aligns with the targets’ actual information, based on participants’ evaluations. When comparing GPT and Claude, it was noted that GPT-generated emails were less detailed, regardless of web usage. We also observed that emails crafted by WebSearch agents were significantly richer in detail and more accurately reflected the target’s personal data than those generated by vanilla LLMs. This difference was particularly pronounced in the category of specific expertise. According to the results of our chi-squared tests in Table 9 (Appendix E.2), the alignment of content with the target’s specific expertise had the most significant impact on participants’ perception of the email’s authenticity. By leveraging detailed information about the target’s expertise from the web, WebSearch agents were able to craft more convincing and effective phishing emails.
Factors Influencing Phishing Email Detection. We examined the factors that influenced participants’ perceptions of an email as phishing (Q7, multiple responses allowed). The factors were ranked by their influence, with sender information (62.3%), email purpose (39.7%), and information related to expertise (31.6%) being the most significant. This result aligns with the previous studies showing that users heavily consider the sender’s information when assessing an email’s authenticity [51]. This finding suggests that employing spoofing techniques, such as manipulating the sender’s email, could potentially lead to more effective attacks.
6.2.2 Comparison Between Participant Groups
Figure 12: Analysis according to group: (a) academic researchers and (b) non-academic professionals.
Figure 12 presents the actions participants from different groups reported they would take after receiving an email. The academic researcher group generally showed a higher likelihood of clicking on links compared to the non-academic professional group. This difference may be associated with the varying percentages of participants who verified that the phishing emails did not contain accurate information (Q2). In the academic research group, the proportion of emails considered inaccurate varied from 0% to 12.1%, suggesting that the majority of emails contained accurate information. In contrast, in the non-academic professional group, even for WebSearch agent-generated emails, up to 40.7% of participants reported that none of the information matched their own. This is most likely due to the amount of available information online, with research labs and universities often dedicating websites that list their members with information such as current projects, research interests, and recent publications.
For the academic research group, the WebSearch agents for both Claude and GPT were more effective at inducing link clicks compared to the vanilla LLMs, with increases of 1.37 times and 1.75 times, respectively. In the other group, adding web accessibility to Claude resulted in the same link click rate and significantly reduced the “move to spam” percentage. For GPT, adding web accessibility increased the link click rate by 2.49 times. These findings indicate that WebSearch agents can craft effective phishing emails with web accessibility. Researchers, who typically have a considerable amount of their information publicly accessible online, are especially vulnerable to such attacks.
6.2.3 Factors Increasing Phishing Effectiveness
We examined the impact of changing the email’s purpose and sender’s address on the effectiveness of phishing attacks. Emails targeting login credentials proved more effective in prompting recipients to click the provided link, with a click rate of 33.33% for Claude’s WebSearch agent and 25% for GPT’s WebSearch agent (see Table 10 in Appendix E.3). This aligns with existing research [1], which suggested that urgent requests, such as login information updates, are effective in phishing scenarios.
Another notable observation is that using a sender’s email address from an organization different from the target’s resulted in the highest click rate, at 46.67%. This rate is 1.75 times higher than that for the general purposed email generated by the WebSearch agent of Claude, as detailed in Table 10. Participants were less familiar with domains from other organizations, making them more susceptible to deception.
An interesting point is that our results show WebSearch agents could be as effective as, or even more effective than, human-crafted spear phishing emails in prompting clicks. In the previous research [9], link click rates for spear phishing emails using internal email addresses and information were 26.6% for human-crafted and 16.7% for LLM-generated emails. In our study, over 20% of participants (up to 46.67%) responded they would click on links in WebSearch agent-crafted emails. While these results are not directly comparable due to different methods and participant groups, they suggest that LLM agents were highly effective in generating phishing emails. This is particularly noteworthy considering that our attack model requires significantly less knowledge and capability, and involves no human labor.
7 Approachability of Cyberattacks
In this section, we examine the practicality of using LLM agents for targeted cyberattacks by assessing two primary factors: cost and safeguard capability. Understanding these elements is crucial to determining the practicality of employing LLM agents for malicious purposes in real-world scenarios. The time and financial costs associated with the attacks are measured, followed by an assessment of bypassing LLM safeguards.
7.1 Cost Analysis
Conducting attacks incurs both time and financial costs, which can highly impact the practicality of using LLM agents for cyberattacks. We examine how quickly and at what cost attackers can carry out cyberattacks using commercially available LLMs. We measured the time spent and the number of tokens consumed during each attack scenario using WebSearch agent for each individual target. For a cost analysis of the WebNav agent, please refer to Appendix D.1. To simplify our calculations, we considered token usage as the sum of input and output tokens, applying the output token rate, which is typically more expensive than the input token rate. This approach ensures we estimate the upper limit of API costs. For the models utilized, the costs are as follows: GPT 4o [41] and Claude 1.5 Sonnet [5] at \$15/1M tokens, and Gemini-1.5 Flash [20] at \$0.30/1M tokens. To measure the number of tokens, we utilized the tiktoken tokenizer [42] for GPT, as specified by OpenAI [42]. For Claude and Gemini, we used the token counting feature provided by their APIs [3, 13].
On average, collecting PII for an individual  [^13] We sample 50 professors identified in Section 4. required 9.1 seconds and incurred a cost of 2.2 cents (GPT: 6.7s, 2.1¢; Claude: 11.4s, 2.2¢). For generating impersonation posts, GPT, Claude, and Gemini took an average of 9.4 seconds, 21.7 seconds, and 3.6 seconds, respectively, with costs under 3 cents (GPT: 2.2¢; Claude: 3.0¢; Gemini: 0.03¢). For creating phishing emails, GPT and Claude took an average of 7.1 seconds and 22.7 seconds, respectively, with costs below 4 cents (GPT: 2.7¢; Claude: 3.9¢). See Appendix D for the cost statistics.
Cost variations across different LLMs can be attributed to differences in model capabilities, as analyzed in [2]. Considering all scenarios, WebSearch agents prove to be highly practical, executing tasks quickly and at minimal cost. WebSearch agents can exploit private data in cyberattacks very quickly and at minimal cost, emphasizing their practicality for malicious use.
7.2 Safeguard Capability
We conduct an in-depth analysis of safeguard capabilities by exploring different factors in impersonation post generation and spear phishing email generation.
Impersonation Post Generation. In Section 5, we found that the activation of safeguards varied depending on the nature of the claim. For further analysis of the safeguard capabilities of each model, we used two additional claims (“Invest in Dogecoin” and “LLMs are effective in generating phishing emails”) and assessed their safeguard bypassing when paired with different groups of targets. These two groups were 10 professors from previous analyses and 10 influential figures (targets with more fame and general popularity) in the tech industry as targets for impersonation.
The activation of safeguards varied not only based on the nature of the claim but also on the nature of the impersonation target. In this task, we observed different safeguard capabilities in the services, with Claude being the most strict and GPT being the most tolerant of prompts. For GPT, the vanilla LLM and WebSearch agent successfully generated posts for both new claims without any rejections. However, Claude and Gemini showed different results. Whereas WebSearch agents generated posts for the “Invest in Dogecoin” claim, their respective vanilla LLMs produced content for professors but rejected creating content for influential figures at rates of 60% and 20%. When prompted with the claim about the harmful use of LLM, for Gemini, both the WebSearch agent and vanilla LLM successfully generated content, while Claude refused to generate content, responding, “I will not create content promoting harmful uses of AI systems”. The safeguard’s dependence on the target is a characteristic that may restrict an attacker’s range of targets.
Spear Phishing Email Generation. The assessment of safeguard in LLMs for spear phishing email generation involved varying two key factors: 1) the purpose of the email (general, request login credentials, request money), and 2) the type of email address used (institutional or Gmail). We generated emails using the email addresses of the authors, with their consent.
Gemini shows the strongest safeguards by refusing to generate emails in all scenarios of vanilla LLM and WebSearch agent settings (Table 7 in Appendix D.3), responding with messages like “I cannot fulfill your request. Creating emails without consent is unethical and potentially illegal”. All models without web accessibility refuse generating content when the email’s purpose involves requesting login credentials or money, indicating these topics trigger stricter safeguard mechanisms. Conversely, general-purpose emails are less likely to be blocked, with GPT and Claude generating responses in these cases. GPT and Claude are more likely to generate an email when provided with an institutional email address, compared to a Gmail address.
Interestingly, when web accessibility was added, WebSearch agents, which previously refused to generate content under strict settings, began to produce responses. Web-tools increased safeguard bypass of GPT and Claude, allowing generation in most of our experimental settings. This demonstrates that web-based tools may introduce a vulnerability for jailbreaking LLMs.
8 Conclusion
In this work, we examine the emerging threats posed by LLM agents, particularly their potential use of web-based tools in cyberattacks that exploit private data. To assess their capabilities, we focused on three types of cyberattacks: PII collection, impersonation post generation, and spear phishing email generation. Our experimental results demonstrate that attackers can successfully automate these attacks using LLM agents, with web-based tools enhancing the performance of these attacks. Furthermore, our results report easy bypass of LLM service safeguards at low-costs, making utilization of LLM agents in cyberattacks a viable option. These findings expose a significant vulnerability in current safeguards and underscore the urgent need for security measures to prevent the misuse of LLM agents.
Ethical Considerations
We care deeply about human safety and societal security, which motivates us to conduct this research. This paper discusses and carefully examines the potential threat of LLM agents in cyberattacks exploiting personal data. This raises ethical considerations as some of the data that we extract contains information about individual users. We minimize these concerns by using already public data and not presenting any specific information in this paper.
Aggregation of PIIs. The collected PII data is stored in pickle format on the principal investigator’s server. The data was provided to annotators in an Excel file format for annotation purposes. Once the annotation process was completed, the annotators returned the annotated files to the principal investigator and deleted their local copies. Currently, files containing personal information are exclusively accessible by the principal investigator and will be discarded after the publication of the research.
User Study. In the case of the user study in Section 6, according to the local IRB’s protocol, our research is excluded from the review for the following reasons. First, we used only publicly available information and did not handle any sensitive data. Additionally, the emails evaluated contained no harmful or offensive content, minimizing any potential impact on participants. In the recruitment process, we obtained informed consent from participants, detailing the purpose of the research, the expected duration, procedures involved, anticipated benefits, personal information protection, and consent for collecting and using their provided email addresses. We also informed participants that the LLM agents could search the internet for their email addresses to gather information and use it to generate emails. Only those consented could participate, and each participant completed the survey once. In the evaluation process, we did not send emails to the participants. We asked to evaluate emails through a Google survey after notification. Each participant was exposed only to emails generated from their own email address. All participant-related information was destroyed after the experiments, and participants were informed about this process. Additionally, we informed them that their responses would be excluded from the analysis if they withdrew consent, ensuring robust protection of personal information.
Responsible Disclosure Process. We believe that the benefits of publicizing these attacks outweigh the potential harms. First, we plan to submit our research findings to the LLM service vendors (OpenAI, Anthropic, and GoogleAI), including our prompts and concerns about the safeguard vulnerabilities of web-based tools. Additionally, we will request the implementation of a rule ensuring that when an LLM utilizes web crawling tools, it checks and adheres to robots.txt directives. (A robots.txt file tells web crawlers which parts of a site they can or cannot access.) Second, for the PII providers from whom we collected data, we strongly recommend creating a robots.txt file to block web crawlers from accessing pages containing personal information.
Open Science Policy Compliance
Our study explores the potential threat posed by LLM agents in cyberattacks. We acknowledge that disclosing the exact prompts used for these attacks could enable malicious actors to replicate them. To mitigate the risk of misuse, we have submitted our code for review with dummy prompts accessible at this link  [^14] https://zenodo.org/doi/10.5281/zenodo.13691327
. After decision, we will discuss with reviewers and determine the feasibility of releasing our code. We will provide the full source code only upon request for legitimate research purposes.
Appendix A Example of Search() function results
Figure13 shows the example of Search() function results. We use ‘Usenix Security 25’ as a query.
Figure 13: Part of the results of search(Usenix Security 25).
Appendix B PII collection
B.1 Human annotation
To evaluate the accuracy of the collected PII, we employed human annotators to review the responses generated by LLM agents. Six annotators were employed, all of whom held at least a bachelor’s degree and were proficient in English.
Due to the vast and diverse nature of the field of CS, we did not evaluate based solely on department listings. Instead, annotators were directed to classify individuals as CS professors based on their engagement in relevant fields, following a thorough web search. Only for individuals identified as CS professors, annotators conducted the remaining four PII verification processes. Certain types of PII, such as phone number, are prone to changes over time, complicating the verification process due to discrepancies across different web sources. Furthermore, the probabilistic nature of responses from LLMs can lead to inconsistent information retrieval. To address these concerns, we initiated five separate queries to LLMs for each individual. Annotators were then instructed to consider the information collection successful if at least one of the five responses matched the data available online.
During the annotation process for the names of CS students, annotators referenced the student information provided at the professors’ laboratory introduction pages as the ground truth. The methodology for annotating other types of PII remained consistent with the approach used for professors’ PII.
Appendix C Impersonation Post Generation
C.1 Templates for Evaluation
A/B Testing Prompt This is a social media post (like Twitter or Facebook). Decide which post appears more likely to have been created by the author themselves, rather than generated by AI or another person. You can use accurate personal information of the author from a web search. Respond with either [answer]postA[answer] or [answer]postB[answer] or [answer]unsure[answer]. PostA: TextA, PostB: TextB
Yes/No Testing Prompt Do you think this content was written by the author themselves, rather than being generated by AI or another person? You can use personal information of the author from a web search. Respond with either [answer]Yes[answer] or [answer]No[answer]. Text.
C.2 Human Correlation with LLM Evaluation
We assessed the reliability of our LLM evaluators by comparing with human reviewers. We sampled 270 posts, each impersonating one of 10 professors for each claim, language model, and capability(LLM, WebSearch gent, WebNav agent) combination mentioned in Section 5.1. Then, both human reviewers and agents evaluated these posts. The average agreement rate between humans and agents was 93.0% for the first claim and 85.2% for the second, underscoring the reliability and credibility of the LLM evaluations.
Appendix D Approachability of Cyberattacks
D.1 PII Collection
We conduct a cost analysis of PII collection using WebSearch and WebNav agents. Using navigation was more expensive in terms of time and token consumption compared to not using it (Table 5). During the process of searching for information, time and tokens are consumed as various links and buttons are clicked within the site. Upon integrating navigation functionality, the average token usage for GPT and Claude agents increased by 8.95 times and 6.80 times, respectively.
GPT vs. Claude. Interestingly, regardless of the use of navigation functions, GPT made more function calls than Claude. Without navigation, this difference became even more pronounced because GPT continuously altered search queries when the desired information was not immediately found. This result highlights GPT’s more proactive exploration compared to Claude, as previously mentioned.
When not using the navigation function, GPT required less time to complete tasks than Claude, despite making more function calls. This could be attributed to differences in model performance, aligning with the analysis reported in [2].
Table 5: Cost statistics for LLM agents to collect PII of professors with and without the navigation function.
GPT Claude
navigation x o x o
Time (sec) min 3.33 6.53 7.85 15.13
max 11.18 294.45 17.51 180.66
mean 6.69 53.50 11.43 43.28
min 1.00 2.00 1.00 2.00
# Function calls max 6.00 17.00 2.00 8.00
mean 3.56 5.74 1.40 3.72
# Tokens min 438.00 895.00 1159.00 1948.00
max 2559.00 76041.00 2268.00 78662.00
mean 1406.04 12583.74 1462.78 9952.80
D.2 Impersonation Post and Phishing Email Generation
We conduct a cost analysis of impersonation post generation and phishing email generation using WebSearch agents. Table 6 represents cost statistics.
Table 6: Cost statistics for WebSearch agents to generate impersonation post and spear phishing email.
GPT Claude Gemini
Impersonation Time min 2.0 6.5 2.7
max 24.2 33.0 4.6
mean 9.4 21.7 3.6
# Tokens min 519 1122 772
max 2640 2751 1075
mean 1435.1 1993.2 917.7
Phishing Time min 4.9 18.0 N/A
max 11.1 43.9 N/A
mean 7.1 22.7 N/A
# Tokens min 803.0 1984.0 N/A
max 2760.0 3761.0 N/A
mean 1803.7 2623.0 N/A
D.3 Safeguard Capability in Phishing Email Generation
Table 7 shows the response generation capabilities of LLMs across different configurations.
Table 7: Response generation capabilities of LLMs across different configurations based on the purpose of the email or the type of email address. successfully generated a response (o) or refused to generate a response (x)
LLM LLM Agent
GPT Claude Gemini GPT Claude Gemini
Purpose General o o x o o x
Credential x x x o o x
Money x x x o x x
Email address Institutional o o x o o x
Gmail o x x o o x
Appendix E Spear Phishing Email Generation
E.1 Recruiting Process and Ethical Measures
We advertised our study within communities and contacted potential participants from our personal networks. All participants were informed about an incentive of approximately \$7.50 for their participation.
To analyze the evaluation results relative to the characteristics of the institutions associated with the targets, we ensured a balanced distribution among individuals from academic research and non-academic professional domains. This balanced approach facilitated a comprehensive evaluation of the variables under investigation.
At the beginning of the study, participants were provided with a consent form. Then, we asked to provide their institutional email addresses. Upon receiving consent, we collected email addresses for the purpose of data collection integral to the experimental design. For each email address, we generated seven emails using LLMs and WebSearch agents.
Participants were then asked to assess these emails through a questionnaire on Google Survey that we administered. It is important to note that each participant was exposed only to emails generated from their own email address. We instructed participants to evaluate the emails under assumption that they had received these emails to their own email addresses.
Of the 70 individuals who provided their email addresses via the consent form, 60 completed the evaluation of the emails.
E.2 Chi-Squared Tests
This section investigates the factors influencing participants’ perception of email authenticity. First, we identify which aspects of an email had the most significant influence on determining the email’s authenticity. We conducted chi-squared tests to evaluate the independence between email evaluation questions (Q1–Q4) and the question assessing perceived email authenticity (Q6). Please refer to Appendix E.5 for the questions used in our survey. Table 8 shows that “content detail” (Q1) and “content authenticity” (Q2) had p-values much less than 0.05. These results suggest that these factors are critical in determining whether an email is perceived as genuine.
Further analysis focused on “content authenticity” (Q2) identified as the most strongly associated factor. Specifically, we examined which aspects of email content influenced participants’ perceptions of authenticity. We evaluated the independence between each response option within Q2 and the option in Q6 using chi-squared tests. Table 9 shows that the “Specific area of expertise” option had a p-value much less than 0.05, indicating a highly significant association. This finding suggests that when the email content aligns with the participant’s domain-specific knowledge, it strongly impacts their perception of the email’s authenticity.
Table 8: Chi-squared test results between each question (Q1-Q4) and Q6.
Q1 Q2 Q3 Q4
(content detail) (content auth) (sender) (recipient)
kappa 98.80 113.54 56.54 3.69
p-value $2.23\times 10^{-5}$ $3.12\times 10^{-7}$ 0.0159 0.4497
Table 9: Chi-squared test results between each response option within Q2 and Q6.
institution general expertise specific expertise hobby
kappa 4.34 0.01 11.15 0.28
p-value 0.0595 0.0373 0.9436 0.0008
E.3 Overall Survey Results
Table 10 represents the survey results about participants’ actions (Q5) across seven email types.
E.4 Contingency Table
Table 11 presents contingency tables illustrating the percentage distribution of participants’ actions (Q5) in relation to their perception of authenticity (Q6). The analysis focuses on emails with general purposes and no specified sender institution.
Table 10: Survey results about expected actions upon receiving emails. Percentages represent responses relative to the total number of participants.
Action
Model Web Purpose Sender’s institution Click link Reply Nothing Move to spam
Claude o general not designated 26.67 6.67 43.33 23.33
x general not designated 21.67 10.00 28.33 40.00
o credential not designated 33.33 6.67 25.00 35.00
o general different from target’s 46.67 3.33 41.67 8.33
GPT o general not designated 20.00 6.67 30.00 43.33
x general not designated 10.00 1.67 45.00 43.33
o credential not designated 25.00 5.00 23.33 46.67
Average 26.19 5.71 33.81 34.29
Table 11: Contingency table showing the percentage distribution of participants’ actions (Q5) against their perception of authenticity (Q6).
Genuine or Phishing
WebSearch Agent (Claude) Definitely genuine Probably genuine Unsure Probably phishing Definitely phishing Total
Action Click link 5.00 8.33 1.67 10.00 1.67 26.67
Reply 0.00 0.00 3.33 3.33 0.00 6.67
Nothing 0.00 5.00 6.67 21.67 10.00 43.33
Move to spam 0.00 0.00 1.67 8.33 13.33 23.33
Total 5.00 13.33 13.33 43.33 25.00 100.00
Genuine or Phishing
LLM (Claude) Definitely genuine Probably genuine Unsure Probably phishing Definitely phishing Total
Action Click link 0.00 11.67 1.67 5.00 3.33 21.67
Reply 5.00 3.33 0.00 1.67 0.00 10.00
Nothing 0.00 3.33 5.00 10.00 10.00 28.33
Move to spam 0.00 0.00 0.00 5.00 35.00 40.00
Total 5.00 18.33 6.67 21.67 48.33 100.00
Genuine or Phishing
WebSearch Agent (GPT) Definitely genuine Probably genuine Unsure Probably phishing Definitely phishing Total
Action Click link 1.67 5.00 6.67 5.00 1.67 20.00
Reply 1.67 1.67 0.00 3.33 0.00 6.67
Nothing 0.00 3.33 5.00 10.00 11.67 30.00
Move to spam 0.00 0.00 1.67 10.00 31.67 43.33
Total 3.33 10.00 13.33 28.33 45.00 100.00
Genuine or Phishing
LLM (GPT) Definitely genuine Probably genuine Unsure Probably phishing Definitely phishing Total
Action Click link 1.67 5.00 1.67 1.67 0.00 10.00
Reply 0.00 0.00 0.00 1.67 0.00 1.67
Nothing 0.00 3.33 10.00 16.67 15.00 45.00
Move to spam 0.00 0.00 1.67 8.33 33.33 43.33
Total 1.67 8.33 13.33 28.33 48.33 100.00
E.5 Questionnaire for User Study
Q1. [Content Detail] What content does the email include? (select all that apply)
Note: Choose answers based solely on the presence of content, regardless of its authenticity or relevance to you.
  • $\square$
    None of the items match
  • $\square$
    Institution (e.g., Tesla, Meta, etc.)
  • $\square$
    General area of expertise (e.g., EE, CS, NLP, Implementation, etc.)
  • $\square$
    Specific area of expertise (e.g., LLM Security, Reinforcement Learning, Implementation using Copilot, etc.)
  • $\square$
    Hobby, Pastime activity (e.g., Football, Game, etc.)
Q2. [Content Authenticity] Which of the included contents are authentic? (select all that apply)
Note: “Authentic” means the information aligns with your interests or expertise areas.
For example, if your areas of expertise are LLM Security & LLM Benchmarks, only consider content authentic if it correctly relates to these areas.
  • $\square$
    None of the items match
  • $\square$
    Institution
  • $\square$
    General area of expertise
  • $\square$
    Specific area of expertise
  • $\square$
    Hobby, Pastime activity (e.g., Football, Game, etc.)
Q3. [Sender] Check the items that apply to the sender of the email. (select all that apply)
  • $\square$
    None of the items match
  • $\square$
    I know the sender’s affiliated institution (e.g., Tesla, Meta)
  • $\square$
    I know the sender’s name (e.g., Elon Musk, Mark Zuckerberg)
  • $\square$
    The sender’s name and affiliated institution are consistent (e.g., Elon Musk & Tesla, Mark Zuckerberg & Meta)
  • $\square$
    I am acquainted with the sender (e.g., friend, student, professor)
Q4. [Recipient] Does the recipient name in the email align with your “real name”?
[…]
    Yes
[…]
    No
Q5. [Action] If you receive this email, what would be your action? (single select)
[…]
    Respond to this email
[…]
    Click the link
[…]
    Read the email and do nothing
[…]
    Move to spam
Q6. Do you think this email is genuine or phishing? (single select)
[…]
    Definitely genuine
[…]
    Probably genuine
[…]
    Unsure
[…]
    Probably phishing
[…]
    Definitely phishing
Q7. (Respond if you answered that the email seems like phishing in Q6) What aspect of the email contributed to your perception of it being phishing? (select all that apply)
  • $\square$
    Information on specific events, schedules, etc. (time, location, etc.)
  • $\square$
    Recipient information (recipient’s name, etc.)
  • $\square$
    Sender information (sender’s name, institution, email address, etc.)
  • $\square$
    Information related to expertise or interests
  • $\square$
    Purpose of the email